OpenStudy (anonymous):

Find the range of he inverse trigonometric function f(x)=tan^-1 (x)

OpenStudy (anonymous):

you don't "find the range" you "define the range"

OpenStudy (anonymous):

by convention the range of arcangent is \((-\frac{\pi}{2},\frac{\pi}{2})\)

OpenStudy (anonymous):

how would you find that answer

OpenStudy (anonymous):

by looking in a book really i am not being silly tangent is periodic, so it is certainly not one to one you can make the inverse into a function by restricting the range of the inverse, which is the same as restricting the domain of tangent by convention the restriction is \((-\frac{\pi}{2},\frac{\pi}{2})\) that is just the convention, which is why you don't "find" it

OpenStudy (anonymous):

okay i understand, so how would one define the range of he inverse trigonometric function f(x)=sin^-1 (x)

OpenStudy (mertsj):

Same way...look in a book to see what mathematicians have decided to define as the range.

OpenStudy (anonymous):

by convention it is \([-\frac{\pi}{2}, \frac{\pi}{2}]\) you need only look in a text, as @Mertsj said

OpenStudy (anonymous):

thanks :)